THE ν_{2} BENDING POTENTIAL OF THE X̃ ^{2}Σ^{+} STATE OF MgNC

\underline{MASARU FUKUSHIMA} and TAKASHI ISHIWATA, Faculty of Information Sciences, Hiroshima City University, Asa-Minami, Hiroshima 731-3194, Japan.

We have generated MgNC in supersonic free jet expansions, and observed the laser induced fluorescence ( LIF ) of the Ã 2Π – X̃ ^{2}Σ^{+} transition. We measured the LIF dispersed fluorescence spectra from the single vibronic levels ( SVL ) of the Mg-N-C bending ( ν_{2} ) mode in the upper Ã ^{2}Π state. The vibronic structures in the SVL dispersed spectra were analyzed according to customary procedure. On the basis of the results, we discuss the vibrational structure of the ground X̃ ^{2}Σ^{+} state of MgNC and the potential surface of the isomerization reaction, MgNC ↔ MgCN. In our previous work^{a}, we also applied the internal rotation model to analyze the bending vibrational structure, and derived the isomerization reaction potential surface, a part of which corresponds to the ν_{2} bending potential. In the internal rotation model, the wavefunctions of the ν_{2} bending mode are expanded using those of the plane free rotor; i.e. this means that we treated only the l = 0 component. On the present work, we expand the wavefunctions using spherical harmonics, and derive the potential using all of the observed bending levels.

aM. Fukushima and T. Ishiwata, 60th International Symposium on Molecular Spectroscopy, TC03 (2005).